# The Regular property of Invariant Rings over a Dedekind Domain

### v1

@@ Abstract.
Let AA be a Dedekind domain and let $K$ be its field of fractions. Let $G\subseteq GL_{n}(A)$ be a finite group. Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing AA). Assume that $|G|$ is invertible in AA. We prove that if $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@
invariant rings, ring of integers of global fields, regular rings, group cohomology.
@@ 1. Introduction
Consider the hypothesis below.
Hypothesis $1.$
Let AA be a Dedekind domain and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing AA). Assume that $|G|$ is invertible in AA.
In this article we establish the following result which is a variant of Theorem 7.2.1 in [1] for polynomial ring over a Dedekind domain.
@@
Assume Hypothesis $1.$ Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Example 1.3.
The conclusion of Theorem 1.2 is true in the following cases.
  1. (1)
    AA is the ring of integers of a global function field i.e. a finite extension of $\mathbb{F}_{q}$(t) where $q$ is some power of a prime $p$, and $p\nmid|G|$.
  2. (2)
    AA is localization at a prime of the ring of integers of a number field and $|G|$ is non-zero in the residue field.
Remark 1.3.1.
We might wonder whether Theorem 1.2 holds if AA is a regular domain of higher dimension. However our technique fails when $\dim A\geq 2$, see 2.3.1.
2. Auxiliary Results
@@
Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is a field). Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$ with $|G|=r$ invertible in $\mathcal{O}/\pi\mathcal{O}$.
@@
So $YU=0$. The image of $U$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is rIrI which is invertible in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ since $r$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Consider the map $U:\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $\mathcal{O}^{n}$ is noetherian module over $\mathcal{O}$, the map is an isomorphism. Thus $U\in GL_{n}(\mathcal{O})$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
@@
Remark 2.3.1.
Our proof of Lemma 2.3 fails when $\mathcal{O}$ has dimension $\geq 2$. Suppose
$\mathcal{O}=k[X,Y]_{(X,Y)}$. Let $v=e_{1}/x^{2}+e_{2}/y^{2}\in K^{2}$ be such that $\sigma(v)=\lambda v$. Then note that for no choice of $t\in\mathcal{O}$ does $tv\in\mathcal{O}^{2}\setminus\mathfrak{m}\mathcal{O}^{2}$. So our induction fails at the base case $n=2$.
3. Proof of Theorem 1.2
Firstly we observe the following.
Hypothesis 3.1.
Consider the case in Hypothesis $1.$ Let $P$ be a maximal ideal of AA. Set $(\mathcal{O},(\pi))=(A_{P},PA_{P})$. Now $(\mathcal{O},(\pi))$ is a DVR and $\mathcal{O}_{\pi}=K$. Also, as $G$ is a subgroup of $GL_{n}(A)$, it is naturally a subgroup of $GL_{n}(\mathcal{O})$. Further since $|G|$ is invertible in AA, we have $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Thus for proving Theorem 1.2, it suffices to prove that $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular. Now let $R=\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$.
Now we give the proof of Theorem 1.2 by proving $S$ is regular in Hypothesis 3.1.
@@
Consider the short exact sequence
@@
Let $K=ker(\pi)=Img(\delta)$ in the above sequence. Now we have that $K$ is $R^{G}/\pi R^{G}$-module. By Theorem 6.5.8 in [3] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. Hence $r\cdot K=0$. As $r$ is a unit in $R^{G}/\pi R^{G}$, we have $K=0$. Hence we have the short exact sequence
@@
Thus $R^{G}/\pi R^{G}\cong(R/\pi R)^{G}$. Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field. Also by Corollary 2.3.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by psuedo-reflections. Thus by Theorem 7.2.1 in [1], $(R/\pi R)^{G}$ is regular. Let $\mathfrak{m}$ be the maximal homogenous ideal of $R^{G}$. Since $\pi\not\in\mathfrak{m}^{2}$. Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎

### v2

# The Regular property of Invariant Rings over Regular Domains

@@ Abstract.
The main result of this paper is a generalization of the theorem of Chevalley-Shephard-Todd to the rings of invariants of pseudo-reflection groups over regular domains. More precisely, let $A$ be a regular domain and let $K$ be its field of fractions. Let $G\subseteq GL_{n}(A)$ be a finite group. Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$). Assume that $|G|$ is invertible in $A$. We prove that if $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@
invariant rings, regular domains, regular rings, group cohomology.
@@ 1. Introduction
We have the famous Chevalley-Shephard-Todd Theorem (Theorem 7.2.1 in [1]) which states:
@@
Let $K$ be a field. Consider a finite group $G\subseteq GL_{n}(K)$. Consider the ring
$K[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $K$). Assume that $|G|$ is invertible in $K$. We have $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections if and only if $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over $K$.
Mundelius had generalized Theorem 1.1 to the rings of invariants of pseudo-reflection groups over Dedekind domains (Theorem 3.8 in [2]).
In this paper we establish the following result which is a further generalization of Theorem 1.1 to the rings of invariants of pseudo-reflection groups over regular domains.
Theorem 1.2.
Let $A$ be a regular domain and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$. If $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections, then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
In Section 2, we prove Theorem 1.2 for the case when $A$ is a DVR as Theorem 2.2. This also leads to an alternate proof for the case of Dedekind domains considered by Mundelius in [2]. In Section 3, we prove Theorem 1.2 for the case when $A$ is a regular local ring as Theorem 3.2. Finally in Section 4, we complete the proof of Theorem 1.2.
2. Proof of Theorem 1.2 for DVRs
@@
Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is the field of fractions of $\mathcal{O}$). Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$. Consider the ring $\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $\mathcal{O}$). Assume that $|G|$ is invertible in $\mathcal{O}$ i.e. $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
In this section, we will prove the following.
Theorem 2.2.
Assume Hypothesis 2.1. Consider $G\subseteq GL_{n}(\mathcal{O}_{\pi})$ is generated by psuedo-reflections. Then $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
@@
So $YU=0$. The image of $U$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is $rI$ which is invertible in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ since $r$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Consider the map $U:\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $\mathcal{O}^{n}$ is noetherian module over $\mathcal{O}$, the map is an isomorphism. Thus $U\in GL_{n}(\mathcal{O})$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
@@
Now we give the proof of Theorem 2.2.
@@
Let $R=\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
@@
By Theorem 6.5.8 in [4] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. As $r$ is a unit in $\mathcal{O}=(R^{G})_{0}$, we have that it is a unit in $R^{G}$. Thus $H^{1}(G,R)=0$. Hence we have the short exact sequence
@@
Thus $R^{G}/\pi R^{G}\cong(R/\pi R)^{G}$. Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field. Also by Corollary 2.4.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by psuedo-reflections. Thus by Theorem 1.1, $(R/\pi R)^{G}$ is regular. Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$. Since $\pi\not\in\mathfrak{M}^{2}$. Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
3. Proof of Theorem 1.2 for regular local rings
Consider the hypothesis below.
Hypothesis 3.1.
Let $(A,\mathfrak{m})$ be a regular local ring and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$.
In this section, we will prove the following.
Theorem 3.2.
Assume Hypothesis 3.1. Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Now consider the following.
Hypothesis 3.3.
Consider the case in Hypothesis 3.1. Let $d=\dim(A)$. Now Theorem 3.2 for the case $d=0$ (i.e. $A$ is a field) follows from Theorem 1.1. Thus we can assume $d\geq 1$. Let $\mathfrak{m}=(x_{1},\dots,x_{d})$. Set $x=x_{1}$ and let $\mathfrak{p}$ be the prime ideal $(x)$.
We have an analogous result to Proposition 2.3 for regular local rings. Essentially the same proof goes through but we provide it for the sake of completeness.
Lemma 3.4.
Assume Hypothesis 3.3. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective.
Proof.
We consider elements of $GL_{n}(A)$ and $GL_{n}(A/\mathfrak{p})$ as matrices. Let $W\in G$ with $\eta(W)=I$. Then $W=I+Y$ where all entries of $Y$ are in $\mathfrak{p}$. Now $(I+Y)^{r}=W^{r}=I$ where $r=|G|$. Thus
$rY+{r\choose 2}Y^{2}+\dots+\ {r\choose{r-1}}Y^{r-1}+Y^{r}=0.$
Set
$U=rI+\ {r\choose 2}Y+\dots+\ {r\choose{r-1}}Y^{r-2}+Y^{r-1}.$
So $YU=0$. The image of $U$ in $GL_{n}(A/\mathfrak{p})$ is $rI$ which is invertible in $GL_{n}(A/\mathfrak{p})$ since $r$ is invertible in $A/\mathfrak{p}$. Consider the map $U:A^{n}\rightarrow A^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $A^{n}$ is noetherian module over $A$, the map is an isomorphism. Thus $U\in GL_{n}(A)$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
Now we give the proof of Theorem 3.2 with notations as in Hypothesis 3.3.
Proof.
We prove the result by induction on $d\geq 1$. Now the result for the case $d=1$ (i.e. $A$ is a DVR) follows from Theorem 2.2. So assume $d\geq 2$ and that the result is known for all regular local rings of dimension d1d-1.
As $A$ is a regular local ring we have that $A_{\mathfrak{p}}$ is a regular local ring of dimension $1$, i.e. a DVR which has $K$ as its field of fractions. Let $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ which is the field of fractions of $A/\mathfrak{p}$. Now $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Thus by Corollary 2.4.1, we have that the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections. Note that we have a commutative diagram of ring homomorphisms,
[…]
where $\mu$ is the standard ring homomorphism of $A/\mathfrak{p}$ to its field of fractions which is injective. As we have the above commutative diagram, we also have the below commutative diagram of group homomorphisms with $\mu$ being injective.
$GL_{n}(A)$$GL_{n}(A/\mathfrak{p})$$GL_{n}(A_{\mathfrak{p}})$$GL_{n}(\kappa(\mathfrak{p}))$$\scriptstyle{\iota}$$\scriptstyle{\gamma}$$\scriptstyle{\mu}$$\scriptstyle{\gamma^{\prime}}$
We have that $\gamma|_{G}:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective by Lemma 3.4. Thus we have that $\mu\circ\gamma|_{G}$ is injective. Now $\gamma^{\prime}\circ\iota|_{G}=\mu\circ\gamma|_{G}$. So the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections. By induction we have $(A/\mathfrak{p}\ [X_{1},X_{2},\dots X_{n}])^{G}$ to be regular. Let $R=A[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
$0\rightarrow R\xrightarrow{x}R\rightarrow R/\mathfrak{p}R\rightarrow 0.$
We have the long exact sequence
$0\rightarrow R^{G}\xrightarrow{x}R^{G}\rightarrow(R/\mathfrak{p}R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{x}H^{1}(G,R)\rightarrow H^{1}(G,R/\mathfrak{p}R)\rightarrow\dots$
By Theorem 6.5.8 in [4] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. As $r$ is a unit in $A=(R^{G})_{0}$, we have that it is a unit in $R^{G}$. Thus $H^{1}(G,R)=0$. Hence we have the short exact sequence
$0\rightarrow R^{G}\xrightarrow{x}R^{G}\rightarrow(R/\mathfrak{p}R)^{G}\rightarrow 0.$
Thus $R^{G}/\mathfrak{p}R^{G}\cong(R/\mathfrak{p}R)^{G}$ which is regular. Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$. Since $x\not\in\mathfrak{M}^{2}$. Thus $R^{G}/\mathfrak{p}R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
4. Proof of Theorem 1.2 for regular domains
Now we complete the proof of Theorem 1.2.
Proof.
Now $A$ is a regular domain and $K$ is its field of fractions. Let $P$ be a maximal ideal of $A$. So $(A_{P},PA_{P})$ is a regular local ring with $K$ as the field of fractions. Also, as $G$ is a subgroup of $GL_{n}(A)$, it is naturally a subgroup of $GL_{n}(A_{P})$. Further since $|G|$ is invertible in $A$, we have $|G|$ is invertible in $A_{P}$. Thus for proving Theorem 1.2, it suffices to prove that $(A_{P}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular which follows from Theorem 3.2. ∎
Acknowledgements: We would like to thank Prof. Gregor Kemper for making us aware of the work of David Mundelius.